Solve by the elimination method.
3x=3+2y
-4/3x+y=-2/3

Answer by algebrahouse.com(1659)

3x = 3 + 2y -------------> 3x - 2y = 3 {subtracted 2y from each side of top equation}
(-4/3)x + y = -2/3 ------> -4x + 3y = -2 {multiplied bottom equation by 3}

3x - 2y = 3 ------> 12x - 8y = 12 {multiplied top equation by 4}
-4x + 3y = -2-----> -12x + 9y = -6 {multiplied bottom equation by 3}

12x - 8y = 12
-12x + 9y = -6
--------------
y = 6 {added the two equations}

3x = 3 + 2y {original top equation}
3x = 3 + 2(6) {substituted 6, in for y, into top equation}
3x = 3 + 12 {multiplied}
3x = 15 {added}
x = 5 {divided each side by 3}

x = 5 and y = 6

For more help from me, visit: www.algebrahouse.com